force of 2N in tangentially applied to a bicycle wheel for 4 seconds, then removed. If the moment of inertia of the wheel is 3 kg m² , The angular velocity of the wheel after 6 seconds. (radius of the wheel 30 cm) *
1
Expert's answer
2020-12-10T11:03:40-0500

There formula of inertia: I = m × r2; (1)

The formula between force, mass and acceleration is: F = m"\times" a; (2)

The formlua among acceleration, velocity, time and radius, angular velocity is: a = "\frac{v}{t} = \frac{w \times r}{t}" ; (3)


We have: I=3 kg "\times" m2, F = 2 N, t = 6 s, R = 30 sm = 0.3 m.


Firstly, find mass by (1) formula:

m = "\frac{I}{r^2} = \frac{3}{0.3^2} = \frac{100}{3}(kg);"


Then, find a by (2) formula:

"a=\frac{F}{m} = \frac{2}{\frac{100}{3}}=\frac{6}{100} (\frac{m}{s^2})" ;


Then, find angular velocity "w" by (3) formula:

"w=\frac{a \times t}{r}=\frac{\frac{6}{100} \times 6}{0.3}=\frac{36}{30}=\frac{6}{5}=1.2 (\frac{1}{s})" ;